# Fine-Tuning Diffusion Generative Models via Rich Preference Optimization

### v4

@@ 4.3 Data Scaling Analysis and Generalization to Diverse Prompt Distributions
Figure 7: RPO improves DPO performance across data scales, prompt sets, and evaluation metrics. The x-axis indicates the number of Pick-a-Pic training samples used for fine-tuning. The first row presents in-distribution results on Pick-a-Pic, while the second and third rows show out-of-distribution performance on the PartiPrompts and HPSv2 prompt sets, respectively. In all plots, stars indicate the performance of the full Diffusion-DPO model, based on publicly released checkpoints from Wallace et al. (2024), which were trained on nearly 1M offline preference pairs.
@@ 4.4 Direct Abalations on VLM Critics
Figure 8: Radar plot comparing RPO fine-tuning performance using preference data generated by different VLM critics (GPT-4o-1-step, GPT-4o-2-step, RichFeedback, Llava-critic).
@@
Table 2: RPO VLM critic ablation via training on 1.4k pairs of preference data generated by different VLM critics.
@@ 5 Discussions
Our RPO pipeline makes a step towards addressing the challenge of creating high-quality synthetic preferences for generative models post-training. Our procedures are straighforward compliment to existing offline RLHF methods like DPO. In addition, although not explored in this paper, our pipeline can also be catered to online algorithms, such as iterative DPO or RL based methods. We consider this as a promising avenue for future research. Our pipeline also combines the existing progress and improvements in two rapidly growing research directions: VLM and image-editing models, which makes our pipeline pretty robust for future improvement by adopting more capable models as plug-ins.
@@ 6 Conclusion
In this work, we present Rich Preference Optimization (RPO), a method that utilizes rich feedback about the prompt image alignment to improve the curation of synthetic preference pairs for fine-tuning text-to-image diffusion models. After extracting actionable editing instructions from VLMs, we employ ControlNet to modify the images, thereby producing a diverse range of refined samples. The edited images are then combined with the original versions and undergo a relabeling process using a reward model to create a curated set of preference data. By further fine-tuning checkpoint models on this synthetic dataset, we significantly enhance the performance of Diffusion-DPO training and achieve greater data efficiency. Moreover, we believe that our pipeline represents a promising direction and avenue for the data curation of synthetic vision language preference data, one that lies between two fast growing literature-VLMs and Image-editing, showcasing significant potential for future research advancements.
Acknowledgement
The works of Hanyang Zhao, Haoxian Chen, and Wenpin Tang are supported by NSF grant DMS-2113779 and DMS-2206038. Haoxian Chen is supported by the Amazon CAIT fellowship. Tingting Ou is supported by NSF grants CNS-2138834 and EEC-2133516. Wenpin Tang is supported by the Columbia Innovation Hub grant, and the Tang Family Assistant Professorship. The works of Haoxian Chen, Hanyang Zhao, and David Yao are part of a ColumbiaCityU/HK collaborative project that is supported by InnoHK Initiative, The Government of the HKSAR and the AIFT Lab.
@@ A.3 Input Prompts to VLMs for Instruction Generation
Appendix B Qualitative Examples
B.1 Qualitative Examples from Feedback Mechanism–VLM Combinations

### v5

@@ 4.3 Data Scaling and Generalization Analysis
Figure 5: RPO improves DPO performance across data scales, prompt sets, and evaluation metrics. The x-axis indicates the number of Pick-a-Pic training samples used for fine-tuning. The first row presents in-distribution results on Pick-a-Pic, while the second row show out-of-distribution performance on the PartiPrompts prompt set. In all plots, stars indicate the performance of the full Diffusion-DPO model, based on publicly released checkpoints from Wallace et al. (2024), which were trained on nearly 1M offline preference pairs.
@@ 4.4 Direct Abalations on VLM Critics
Table 2: RPO VLM critic ablation via training on 1.4k preference pairs.
@@ 4.5 Effectiveness of Critic-Instruction-Editing Data Curation and RPO Fine-Tuning
Figure 6: Performance comparison of RPO across different VLM critics.
@@ 5 Discussions
Our RPO pipeline makes a step towards addressing the challenge of creating high-quality synthetic preferences for generative models post-training. Our procedures are straightforward compliment to existing offline RLHF methods like DPO. In addition, although not explored in this paper, our pipeline can also be catered to online algorithms, such as iterative DPO or RL based methods. We consider this as a promising avenue for future research. Our pipeline also combines the existing progress and improvements in two rapidly growing research directions: VLM and image-editing models, which makes our pipeline pretty robust for future improvement by adopting more capable models as plug-ins.
@@ 6 Conclusion
In this work, we present Rich Preference Optimization (RPO), a method that utilizes rich feedback about the prompt image alignment from VLMs to improve the curation of synthetic preference pairs for fine-tuning text-to-image diffusion models. After extracting actionable editing instructions from VLMs, we employ ControlNet to modify the images, thereby producing a diverse range of refined samples. The edited images are then combined with the original versions and undergo a relabeling process using a reward model to create a curated set of preference data. By further fine-tuning checkpoint models on this synthetic dataset, we significantly enhance the performance of Diffusion-DPO training and achieve greater data efficiency. Moreover, we believe that our pipeline represents a promising direction and avenue for the data curation of synthetic vision language preference data, one that lies between two fast growing literature-VLMs and Image-editing, showcasing significant potential for future research advancements.
7 Limitations
Given the popularity of our research direction and huge amount of available models, we are only able to test on a set of representative models, which we believe that already provides valuable insights and potentials of our methods even that our tried models are not exclusive. There is also possible risk of image generation models, but our main contribution of this paper is instead novel training recipes/methodologies. In addition, we use AI to assist writing, help with figures design and more efficient coding.
@@ A.3 Input Prompts to VLMs for Instruction Generation
Appendix B ControlNet Examples
Appendix C Qualitative Examples
C.1 Qualitative Comparison of Base Model and Fine-tuned Variants
In this section, we provide a side-by-side comparison of three distinct model configurations: the base SDXL 1.0 , the standard DPO-SDXL-100k baseline , and our RPO-augmented version, DPO-SDXL-100k + RPO100k. As discussed in the main body, while the DPO-SDXL-100k baseline improves general image fidelity, it often fails to coordinate multiple complex prompt elements simultaneously.
As shown in Figure 8, the RPO-enhanced model demonstrates a significantly higher capacity for semantic connectivity. For instance, in the "samurai pizza cat" and "tiger wearing casual outfit" examples, the RPO-enhanced model demonstrates a significantly higher capacity for semantic connectivity, accurately rendering specific clothing and objects that the baseline models omit or distort. These results confirm that rich preference signals provide the nuanced guidance necessary for the model to learn deep relationships between disparate prompt concepts.
C.2 Qualitative Examples from Feedback Mechanism–VLM Combinations
@@
Appendix D Failure Case Analysis
While our RPO pipeline produces informative preference pairs and improves downstream alignment, we also observed several systematic failure cases in data curation that limit the effectiveness of preference data pair curation. To diagnose these issues, we screened examples where the majority of reward models reported negative improvement after editing. Our analysis suggests that most failures originate in the instruction generation stage, with two recurring patterns:
  • Stray content. Generated instructions occasionally introduce content not mentioned in the critique or even contradictory to it. For example, the instruction in Case 1 asked to replace ice cream with a giant chocolate bar, despite the original prompt containing no mention of chocolate.
  • Non-executable instructions. In other cases, the editing instructions require targets absent in the image. For instance, in Case 2, an instruction requested “removing IKEA branding” when no such branding was present.
Beyond these instruction-level errors, we identified two additional sources of failure during the image editing stage:
  • Failure to follow instructions. Even when the instructions are valid, the editing model sometimes diverges from them. For example, in Case 3, an instruction to correct a road sign from “TURN EFT” to “TURN LEFT” was not successfully completed by ControlNet.
  • Low editing quality in human-related images. Edits involving human faces or bodies often introduce artifacts or unrealistic distortions, leading to degraded image fidelity. An example is shown in Case 4.
Failure Case 1: Stray content Prompt: Giant ice cream cone melting and creating a river through a city Critique: 1. Deviations from the Prompt The image depicts a giant ice cream cone melting and creating a river through a city, which aligns well with the prompt. However, there are a few deviations: Scale: The ice cream cone is indeed giant in scale, but it appears to be more of an artistic representation rather than a realistic one. This deviation might be intentional to create a surreal effect. Melting Effect: The melting effect is exaggerated, with the ice cream cone appearing to be melting into a river. This effect is visually striking and aligns with the prompt’s description of “melting and creating a river.” 2. Keywords or Phrases Misinterpreted Giant Ice Cream Cone: The image accurately represents this keyword, showing a massive ice cream cone. Melting: The melting effect is well-executed, though it might be slightly overdone. River Through a City: The image correctly depicts a river flowing through a city, although the city appears to be frozen or covered in snow, which adds an interesting twist. 3. Additional Issues Artifacts and Distortions: There are no significant artifacts or distortions visible in the image. The visual quality is high, with clear details and vibrant colors. Low Detail: The image has a high level of detail, especially in the cityscape surrounding the ice cream cone. The textures and structures of the buildings are well-defined. Conclusion Overall, the image is a creative and imaginative interpretation of the prompt. It captures the essence of the giant ice cream cone melting and creating a river through a city, with only minor deviations that enhance the surreal and artistic nature of the scene. The image is visually striking and effectively conveys the intended concept. Instruction: Replace the ice cream with a giant chocolate bar; change the river to a stream of milk.[Uncaptioned image] Original (left) and Edited (right).
Failure Case 2: Non-executable instructions Prompt: Detailed Ikea ad, messy Room, perfect Lighting and shadows Critique: 1. Deviations from the Prompt Messy Room: The room appears tidy and well-organized, which contradicts the prompt’s request for a “messy” room. Perfect Lighting and Shadows: The lighting in the room is indeed well-balanced and creates natural shadows, but it does not appear overly dramatic or exaggerated as one might expect in a perfect lighting scenario. 2. Poorly Represented or Misinterpreted Keywords Ikea Ad: The image does not clearly represent an Ikea product or advertisement. It shows a room with various items that could be associated with an Ikea style, but there are no explicit signs of being an Ikea ad. Perfect Lighting and Shadows: While the lighting is well-balanced, it does not have the dramatic effect that would make it “perfect” according to the prompt. 3. Additional Issues Artifacts and Distortions: There are no visible artifacts or distortions in the image. Low Detail: The image has a high level of detail, with clear textures and sharpness in the objects and furniture. Conclusion The image does not fully align with the prompt due to the room being tidy and the lighting being well-balanced rather than dramatic. The keywords “Ikea Ad” and “Perfect Lighting and Shadows” are not well-represented in the image. Overall, the image quality is high, but it does not meet the specific criteria set by the prompt. Instruction: Add cluttered furniture; enhance lighting contrast; remove IKEA branding.[Uncaptioned image] Original (left) and Edited (right).
Failure Case 3: Failure to follow instructions Prompt: A road sign with text saying “turn left for food” Critique: 1. Deviations from the Prompt The image provided does not match the prompt accurately. The prompt specifies a road sign with text saying “turn left for food,” while the image shows a road sign that says “turn eft for food.” The word “eft” is incorrect and does not align with the prompt’s instruction to say “left.” 2. Keywords or Phrases Misinterpreted “Turn Left for Food”: The prompt correctly identifies the intended message, but the image incorrectly uses “eft” instead of “left.” “Turn EFT for Food”: The image uses “eft” instead of “left,” which is a deviation from the prompt. 3. Additional Issues Artifacts and Distortions: There are no visible artifacts or distortions in the image. Low Detail: The image has a moderate level of detail, but the main issue is the incorrect wording on the sign, which affects the overall quality and accuracy. Conclusion The image deviates significantly from the prompt due to the incorrect wording on the road sign. This misinterpretation of the prompt’s instructions results in a noticeable discrepancy between the expected and actual content of the image. Instruction: Correct “TURN EFT” to “TURN LEFT”; Change “FOOD” to “RESTAURANT”.[Uncaptioned image] Original (left) and Edited (right).
Failure Case 4: Low editing quality in human-related images Prompt: photo of asian little ballet dancers resting in dance studio, nikon D5 Critique: 1. Deviations from the Prompt Photography Equipment: The image does not show any specific brand or type of camera, such as a Nikon D5, which is mentioned in the prompt. Setting: The image appears to be taken in a dance studio, but the specific details about the studio (e.g., lighting, decor) are not clearly visible or described. Dancers: The image shows four young dancers, which aligns with the prompt’s mention of “little ballet dancers.” 2. Keywords and Phrases Misinterpreted Asian: The prompt specifies “asian,” but the image does not provide any information about the ethnicity of the dancers. Resting: The dancers are sitting and appear to be resting, which aligns with the prompt. 3. Additional Issues Artifacts and Distortions: There are no apparent artifacts or distortions in the image. Detail: The image has a clear focus on the dancers, but the background details are not very distinct. Conclusion The image generally aligns with the prompt regarding the presence of young ballet dancers in a studio setting. However, it misses specific details about the photography equipment and the exact nature of the studio environment. The prompt’s mention of “asian” is not verifiable from the image provided. Overall, the image quality is good, with clear focus on the subjects, but it could benefit from more detailed context. Instruction: Add a floral backdrop behind the dancers; adjust lighting for a softer glow.[Uncaptioned image] Original (left) and Edited (right).
Appendix E Additional Generalization and Scalability Analysis
To complement the analysis in the main paper, we additionally conducted experiments on a second prompt set using the HPSv2 benchmark, together with an extended suite of reward models that place greater emphasis on stylistic attributes. The results in Figure 13 further confirm that the RPO pipeline consistently outperforms the Diffusion-DPO baseline across all training sizes and scales effectively. Moreover, RPO demonstrates stronger efficiency and generalization: with a training size of 100K, RPO outperforms both the Diffusion-DPO model and its checkpoint trained on nearly 1M preference pairs across all the metrics.
Figure 13: Style-specific scaling performance on the HPSv2 dataset. RPO-enhanced models (green) consistently outperform the Diffusion-DPO baseline (blue) and the 1M-sample "Full Diffusion-DPO" baseline (star) across all visual categories.